Question 13 Calculate the number of subsets. {x 1 x is a day of the week) 64 49 127 128
Answer #1

Given,

X = Days of week.

n = number of days in a week = 7

Formula for number of subsets =2^n

2^n=2^7

=128

Number of subsets is 128.

Hence,D)128 is a correct option.